OpenStudy (pphalke):

A home has a triangular backyard. The second angle of the triangle is 7 degrees more than the first angle. The third angle is 13 degrees more than two times the first angle. Find the angles of the triangular yard?

OpenStudy (anonymous):

let the first angle=x second angle=x+7 third angle=2x+13 x+x+7+2x+13=180 ?

OpenStudy (pphalke):

4x+20=180 is that right?

Directrix (directrix):

Yes, it is correct. Continue to solve for x. 4x+20=180

Directrix (directrix):

@pphalke

OpenStudy (pphalke):

is that the answer for the first angle ? what do i do to get the other angles?

OpenStudy (pphalke):

@Directrix

Directrix (directrix):

>>4x+20=180 is the answer to nothing as it stands. You have to solve this for x. Begin by subtracting 20 from both sides of the equation. 4x + 20 = 180 - 20 -20 --------------------- 4x = 160 x = ? What is 160 divided by 4?

OpenStudy (pphalke):

40?

Directrix (directrix):

x = 40, yes Find the measures: let the first angle=x = 40 as we just learned second angle=x+7 = 40 + 7 = third angle=2x+13 = 2*40 + 13 = What are the measures of the second and third angles?